HUBBLE SPACE TELESCOPE DAILY REPORT #2477 PERIOD COVERED: 0000Z (UTC) 10/12/99 - 0000Z (UTC) 10/13/99 Daily Status Report as of 286/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.2 Completed Three Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as planned with no anomalies. 1.3 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.4 Completed WF/PC-2 8243 (SINS: The Supernova Intensive Study - Cycle 8) The WF/PC-2 was used to make observations of supernova LMC-SN1987. Supernovae reveal the last stages of stellar evolution, create the chemical history of the Universe, energize the interstellar gas, and occupy the center of attention for the extragalactic distance scale and cosmology. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/CCD/MA1 8241 (A Snapshot Survey of Interstellar Absorption Lines) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make high resolution Echelle spectrographic observations of the standard ultraviolet star HD62714 as part of a snapshot survey of interstellar absorption lines. A wavelength calibration image was also taken. The observations were executed as scheduled with no anomalies. 1.6 Completed STIS/CCD/MA1 8329 (Double Degenerates Among DAO White Dwarfs) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of the white dwarf star GD561. The evolution of DAO white dwarfs, which have hybrid spectra showing both hydrogen and helium lines, is not well understood. A wavelength calibration image was also taken. The observations were executed as planned with no anomalies. 1.7 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.8 Completed WF/PC-2 8160 (The Optical Counterpart of the Soft Gamma-Ray Repeater 1900+14) The WF/PC-2 was used to make observations of NGC-104 to search for the optical counterpart of the soft gamma-ray repeater 1900+14. There are only four known objects of this type. They are likely an exotic class of neutron star known as magnetars with ultra-strong surface magnetic fields {10^14- 10^15 G}. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 12 Successful: 12 2.2 FHST Updates: Scheduled: 11 Successful: 11 2.3 Operations Notes: The SSR EDAC error counter was cleared three times using ROP SR-1. The NICMOS FPA temperature sensor for detector 1 was enabled at 285/1959Z. ESTR-3 was re-conditioned, as directed by ROP RD-7, at 286/0500Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH